# Daniel San José Pro, Oliver Hausdörfer, Ralf Römer, Maximilian Dösch, Martin Schuck and Angela P. Schöllig

### v1

@@ Abstract
Index Terms:
Compliance and Impedance Control, Telerobotics and Teleoperation, Software-Hardware Integration for Robot Systems Deep Learning in Grasping and Manipulation, Reinforcement Learning, Imitation Learning,
@@ III-A Task Torque
where Log is the logarithmic map used to map a rotation difference to the tangent space of $SO(3)$. This can be seen as an angular velocity that turns the end-effector rotation towards the target rotation [20]. Defining the error in different frames allows the user to define anisotropic stiffness and damping behavior of the controller with respect to the end-effector or base frame.
@@ III-F Target Wrench
Using our implementations, it is possible to command a wrench at the end-effector to apply external forces and torques. The term is computed as
[…]
@@ III-H Additional Features
We provide additional features to ensure safety during operation. First, the user can limit the joint torques as well as their rates of change. To avoid fast movements, targets are filtered with an exponential moving average filter. Further, the tracking error $\mathbf{e}$ can be clipped, which has the advantage of allowing for highly stiff and therefore precise control while not accelerating too fast or in undesired directions if the target pose is far away and the joint torques are limited [14].
For teleoperation setups with a leader and follower robot, we provide controllers with force-torque feedback to aid data collection. If the follower robot can measure its wrenches at the end-effector, a corresponding feedback term can be added to the leader robot: $\bm{\tau}_{\text{fb}}=-k_{\text{p,fb}}\mathbf{J}^{\top}\mathcal{F}_{\text{follower}}-k_{\text{d,fb}}\dot{\bf{q}}_{\text{leader}}$, where $k_{\text{p,fb}}\in\mathbb{R}$ and $k_{\text{d,fb}}\in\mathbb{R}$ are the feedback and damping gains.

### v2

# Daniel San José Pro, Oliver Hausdörfer, Ralf Römer, Maximilian Dösch, Martin Schuck and Angela P. Schoellig

@@ III-A Task Torque
where Log is the logarithmic map used to map a rotation difference to the tangent space of $SO(3)$. This can be seen as an angular velocity that turns the end-effector rotation towards the target rotation [solà2021microlietheorystate]. Defining the error in different frames allows the user to define anisotropic stiffness and damping behavior of the controller with respect to the end-effector or base frame.
@@ III-H Additional Features
We provide additional safety features including joint torque and rate limits, exponential filtering of targets to avoid fast movements, and tracking error clipping for precise yet safe control when targets are distant [22]. For teleoperation, we implement force-torque feedback controllers where follower forces are reflected to the leader as $\bm{\tau}_{\text{fb}}=-k_{\text{p,fb}}\mathbf{J}^{\top}\mathcal{F}_{\text{follower}}-k_{\text{d,fb}}\dot{\bf{q}}_{\text{leader}}$.